# Sampling on infinite-dimensional Paley-Wiener spaces on graphs

### v3

@@ 1 Introduction
The secondary goal of this work is to provide a brief survey on this topic, as the original results of [11] are still being cited in the infinite-dimensional setting without acknowledgment of the erratum (see, e.g., [9], where Section 5 contains a comparison between the sampling set found in that paper and one provided by an example from [11]. Since those results are presented for the square lattice, they should not be taken into account, as their validity had remained unproven until now).
The paper is organized as follows: Section 2.1 is a brief overview on the known sampling results in the continuous case, Section 2.2 contains the main points of Pesenson’s work, along with comments on why they are problematic (Theorem 2.8). Section 3 provides the sampling results for finite and infinite-dimensional Paley-Wiener spaces we are looking for (Theorems 3.1 and 3.2). Section 4 contains a thorough discussion on [6]. First, we will highlight the difference in the methodology (and thus in the results) with respect to our work. Then, we will provide an example of a sampling set which is valid in both cases, and show how our work leads to a sharper constraint, which allows for more Paley-Wiener spaces to be sampled on the same set.
@@ 3 Sampling theorems for infinite-dimensional Paley-Wiener spaces
Theorem 3.2 allows us to consider only the complement of $\lambda$-sets in order to find useful frame-reconstruction sampling sets for a Paley-Wiener space $PW_{\omega}(G)$. However, this result is not enough to establish a full characerization of such subsets, since for a fixed $\omega>0$, we have that $\lambda\omega\leq 1+c_{\omega}$, while Theorem 3.1 gave us the condition $\lambda\omega<1$.
Potential developments of this work might include whether or not the characterization holds whenever $\lambda$ is the smallest possible value for which 2.9 is true. Another open question, under the assumptions on the graph considered in this work, is finding out if there exist uniqueness sets which are not the complement of $\lambda$-sets.
@@ 4.1 Framework
The assumptions on the graph G are the same that have been considered in this work, with the exception of uniformly weighted graphs. This means that the finite-degree assumption can be relaxed: the authors consider a weight function $w:V\times V\longrightarrow\mathbb{R}_{+}$ and they ask that $\sum_{v\in V}w(u,v)<\infty$ for all $u\in V$.
@@
The first essential remark we need to make is that although $\Delta$ is bounded, linear, and positive, it is not equivalent to $\mathcal{L}$ whenever the degree of $G$ is not homogeneous and the graph is not uniformly weighted, with weight $w(u,v)=1/d(G)$ for all $u,v\in V$. So, the results found in [6] do not actually commute with those found here except for the case we just stated. Still, we can consider uniformly weighted graphs such as $\mathbb{Z}^{n}$-lattices or homogeneous trees, equipped with the counting measure and uniform weight $1/d(G)$. We will discuss the differences in these scenarios in the next subsection.
@@ 4.3 Example: comparison on the line $\mathbb{Z}$
Now we perform the due computations to verify which Paley-Wiener spaces can be sampled on this set, starting from the ones provided in [11]:
  • Section 5 from [11] states that every finite sequence of consecutive vertices of lenght N is a $\lambda$-set with
@@
  • Now we just need to compare the two trhesholds. It is easy to show that
@@
    is a non-negative sequence that goes to 0 as $n\rightarrow\infty$ (it is 0 only at $a_{1}=2sin^{2}\left(\frac{\pi}{6}\right)-\frac{1}{2}=0$. It follows that our threshold is sharper for any $k\in 2\mathbb{N}\setminus\{2,4\}$ and it is the same for $k=4$.

### v4

# arXiv:2511.17343v4[math.FA] 02 Mar 2026

@@ 1 Introduction
The secondary goal of this work is to provide a brief note on this topic, as the original results of [12] are still being cited in the infinite-dimensional setting without acknowledgment of the erratum (see, e.g., [10], where Section 5 contains a comparison between the sampling set found in that paper and one provided by an example from [12]. Since those results are presented for the square lattice, they should not be taken into account, as their validity had remained unproven until now).
The paper is organized as follows: Section 2.1 is a brief overview on the known sampling results in the continuous case, Section 2.2 contains the main points of Pesenson’s work, along with comments on why they are problematic (Theorem 2.8). Section 3 provides the sampling results for finite and infinite-dimensional Paley-Wiener spaces we are looking for (Theorems 3.1 and 3.2). Section 4 contains a thorough discussion on [7]. First, we will highlight the difference in the methodology (and thus in the results) with respect to our work. Then, we will provide an example of a sampling set which is valid in both cases, and show how our work leads to a sharper constraint, which allows for more Paley-Wiener spaces to be sampled on the same set. Finally, section 5 will be devoted to finding notable sampling examples on homogeneous trees.
@@ 3 Sampling theorems for infinite-dimensional Paley-Wiener spaces
Theorem 3.2 allows us to consider only the complement of $\lambda$-sets in order to find useful frame-reconstruction sampling sets for a Paley-Wiener space $PW_{\omega}(G)$. However, this result is not enough to establish a full characerization of such subsets, since for a fixed $\omega>0$, we have that $\lambda\omega\leq 1+c_{\omega}$, while Theorem 3.1 gave us the condition $\lambda\omega<1$. In order to dig deeper into this connection and find out whether or not it would be possible to characterize sampling sets via the $\lambda$-constant of their complement, one should start by trying to characerize $c_{\omega}$ in terms of the geometric constraints of the graph. Some work has already been done in this sense, and some results where the sampling threshold depends on geometrical features of the graph have been published (see e.g. [7], [14]). However, since these results are not sharp, they do not provide a sharp $\lambda$-constant for their complements, so the issue still persits. It seems probable that the assumption of finite-geometry (or, in a more general setting, the boundedness of the weight function) would play a crucial role into finding the desired characterisation. This problem is however very difficult to tackle, and it is not precisely on the scope of this work, so we will leave this question open.
@@ 4.1 Framework
The assumptions on the graph G are the same that have been considered in this work, with the exception of uniformly weighted graphs. This means that the finite-degree assumption can be relaxed: the authors consider a non-negative weight function $w:V\times V\longrightarrow\mathbb{R}_{+}$ and they ask that $\sum_{v\in V}w(u,v)<\infty$ for all $u\in V$.
@@
The first essential remark we need to make is that although $\Delta$ is bounded, linear, and positive, it is not equivalent to $\mathcal{L}$ whenever the degree of $G$ is not homogeneous and the graph is not uniformly weighted, with weight $w(u,v)=1/d(G)$ for all $(u,v)$ in $E$. So, the results found in [7] do not actually commute with those found here except for the case we just stated. Still, we can consider uniformly weighted graphs such as $\mathbb{Z}^{n}$-lattices or homogeneous trees, equipped with the counting measure and uniform weight $1/d(G)$. We will discuss the differences in these scenarios in the next subsection.
@@ 4.3 Example: comparison on the line $\mathbb{Z}$
Now we perform the due computations to verify which Paley-Wiener spaces can be sampled on this set, starting from the ones provided by [12]:
  • Lemma 5.3 from [12] states that every finite sequence of consecutive vertices of lenght N is a $\lambda$-set with
@@
  • Now we just need to compare the two thresholds. It is easy to show that
@@
    is a non-negative sequence that goes to 0 as $n\rightarrow\infty$ (it is 0 only at $a_{1}=2sin^{2}\left(\frac{\pi}{6}\right)-\frac{1}{2}=0$). It follows that our threshold is sharper for any $k\in 2\mathbb{N}\setminus\{2,4\}$ and it is the same for $k=4$.
5 Example: spheres on homogeneous trees
In this section, we are going to look into same examples of sampling sets on homogeneous trees that can be found through 3.1. In [12], some examples of this kind are given; however, the reconstruction theorem from [12] was not to be considered valid since it was later corrected in [11] and adapted to finite-dimennsional Paley-Wiener spaces. In light of 3.1 we can now recover those examples and be sure of their validity. Since the examples found in [12] are almost trivial, in the next subsection we will try to find some more notable examples.
5.1 Framework
We consider a homogeneous tree, i.e. a denumerable, connected graph with no cycles and homogeneous degree $d(G)=q+1$, $q\geq 2$. Again, we ask for uniformly weighted edges and we adopt the counting measure on $V$. We fix a vertex of the tree and we refer to it as its “root”. We then indicate with $l_{m}$ the sphere of radius $m$ with respect to the root, with $m\in\mathbb{N}$. We will say that a vertex belongs to $l_{m}$ if there are $m$ edges “separating” it from the root (so the root belongs to $l_{0}$, its neighbours to $l_{1}$, their neighbours to $l_{2}$ and so on).
A very deep and proficient theory has been developed in [4] for these kind of trees. This theory is specialised to radial functions, i.e. $L^{2}$-functions such that they take the same value on $l_{m}$. With a slight abuse of notation, we will denote the space of radial functions on the homogeneous tree $G$ with $L^{2}(G)$. In the same fashion, we will use $PW_{\omega}(G)$ to denote spherical Paley-Wiener spaces defined as
$PW_{\omega}(G):=\left\{f\in L^{2}(G):\mathrm{supp}(Uf)\subseteq[0,\omega]\right\}.$
It is known from [4] that the Laplacian operator on radial functions has spectrum
$\sigma(\mathcal{L})=[1-\rho(q),1+\rho(q)],\qquad\rho(q)=\frac{2q^{1/2}}{q+1}.$ (5.1)
In the next subsection, we are going to exploit this fact together with the properties of radial functions to study sampling sets $W$ of the kind
$W=\bigcup_{m\in\mathbb{N}}l_{km}\cup l_{km+1},\quad k\in\mathbb{N},\quad k>2.$
5.2 Explicit computations for $W$
We will at first compute the explicit $\lambda$-constant for $S:=V\setminus W$ and then explain why we chose these kind of sets. It is immediate to notice that the discrete Laplacian operator on homogeneous graphs can be rewritten in the matrix form:
$\mathcal{L}=I-\frac{1}{d(G)}A,$ (5.2)
where $A$ is the adjacency matrix of $G$, i.e. the symmetric matrix such that $A_{u,v}=1$ if and only if $(u,v)$ belongs to $E$, and it is 0 otherwise. We fix $k\in\mathbb{N}$, $k>2$, and we consider
$W=\bigcup_{m\in\mathbb{N}}l_{km}\cup l_{km+1}.$
We take a fixed $m$ and we denote with $v_{km+1}$ one of the vertices of $l_{km+1}$. We then consider the $q$ sons of $v_{km+1}$ as roots of $q$-ary finite trees of depth k2k-2. This operation is schematised in the following picture.
We denote these kind of finite sub-trees with $T_{j}$. We can now rewrite the discrete Laplacian operator on $T_{j}$ as
$\mathcal{L}_{j}=I_{j}-\frac{1}{q+1}A_{j},$ (5.3)
where $I_{j}$ and $A_{j}$ are the infinite block-matrices such that they are non-zero only in the first $n\times n$ blocks, with $n=\sum_{i=0}^{k-3}q^{i}$. These non-zero blocks correspond respectively to the $n\times n$ identity matrix and to the adjacency matrix of $T_{j}$. This allows us to simplify the calculations for the eigenvalues of $L_{j}$: we just need to know the eigenvalues $A_{j}$. From Theorem 1 in [2], we know that the set of the eigenvalues of the adjacency matrix of a q-ary finite tree of depth k2k-2 is
$\bigcup_{2\leq m\leq k}\left\{2\sqrt{q}cos\left(\frac{l\pi}{m}\right):1\leq l\leq m\right\}.$
It is clear that the maximum eigenvalue of the adjacency matrix is $\mu_{\mathrm{max}}:=2\sqrt{q}cos\left(\frac{\pi}{k}\right)$, so the minimal eigenvalue of $\mathcal{L}_{j}$ is given by
$\eta_{\mathrm{min}}:=1-\frac{1}{q+1}\mu_{max}=1-\frac{2\sqrt{q}}{q+1}cos\left(\frac{\pi}{k}\right)>0.$
This means that, for each $\varphi_{j}$ supported on $T_{j}$, we have
$\|\varphi_{j}\|_{2}\leq\frac{1}{\eta_{\mathrm{min}}}\|\mathcal{L}_{j}\varphi_{j}\|_{2}.$ (5.4)
Now we consider $S=V\setminus W$ and we compute its $\lambda$-constant. Of course, $S=\bigcup_{j\in\mathbb{N}}T_{j}$. Since the $T_{j}$’s are disjointed, for all $\varphi$ in $L^{2}(S)$ we have
$\|\varphi\|^{2}_{2}=\sum_{j\in\mathbb{N}}\|\varphi_{j}\|^{2}_{2}\leq\sum_{j\in\mathbb{N}}\frac{1}{\eta_{\mathrm{min}}^{2}}\|\mathcal{L}_{j}\varphi_{j}\|^{2}_{2}=\frac{1}{\eta_{\mathrm{min}}^{2}}\sum_{j\in\mathbb{N}}\|\mathcal{L}_{j}\varphi_{j}\|^{2}_{2}.$
Now we have to estimate $\sum_{j\in\mathbb{N}}\|\mathcal{L}_{j}\varphi_{j}\|^{2}_{2}$ in terms of $\|\mathcal{L}\varphi\|^{2}_{2}$. This time, since the $T_{j}$’s closures are not pairwise disjoint, we cannot appeal to Lemma 2.7. However, we can exploit the radial simmetry of $\varphi$: If we fix a vertex on $l_{km+1}$, and we denote it with $v_{km+1}$, we have that
$|\mathcal{L}_{j}\varphi_{j}(v_{km+1})|=\frac{1}{(q+1)^{2}}|\varphi(v_{km+2})|,\qquad v_{km+2}\in l_{km+2}.$
This means that
$\sum_{j\in\mathbb{N}}|\mathcal{L}_{j}\varphi_{j}(v_{km+1})|^{2}=\frac{q}{(q+1)^{2}}|\varphi(v_{km+2})|^{2}=\frac{1}{q}|\mathcal{L}\varphi(v_{km+1})|^{2},$
moreover, for any $v$ in $V\setminus\bigcup_{m\in\mathbb{N}}l_{km+1}$, we have
$\sum_{j\in\mathbb{N}}|\mathcal{L}_{j}\varphi_{j}(v)|^{2}=|\mathcal{L}\varphi(v)|^{2}.$
Thus $\sum_{j\in\mathbb{N}}\|\mathcal{L}_{j}\varphi_{j}\|^{2}_{2}\leq\|\mathcal{L}\varphi\|^{2}_{2}$. In the end we get
$\|\varphi\|_{2}\leq\frac{1}{\eta_{\mathrm{min}}}\|\mathcal{L}\varphi\|_{2},\qquad\varphi\in L^{2}(S).$ (5.5)
By Theorem 3.1, we have that $W$ is a sampling set for all radial Paley-Wiener spaces $PW_{\omega}(G)$ such that
$\omega<\eta_{\mathrm{min}}.$
Since the spectrum of $\mathcal{L}$ is the interval given by 5.1, the sampling range is given by
$\omega\in\left[1-\frac{2\sqrt{q}}{q+1},1-\frac{2\sqrt{q}}{q+1}cos\left(\frac{\pi}{k}\right)\right).$
As one might expect, as $k$ increases (i.e. the distance between elements of the sampling set increases), the sampling range gets narrower.
5.3 Final remarks on the choice of the sampling set
The choice of a samplig set built up by layers of two consecutive spheres was motivated by two main reasons (other than trying to make it as sparse as possible): first of all, since we are dealing with radial functions, it makes sense to sample on the whole sphere, since such functions take the same value at each $l_{m}$. Second of all, if we only considered the sampling set made up by separated spheres $\bigcup_{m\in\mathbb{N}}l_{km}$, $k>1$, we would have not been able to estimate $\sum_{j\in\mathbb{N}}\|\mathcal{L}_{j}\varphi_{j}\|_{2}$ in terms of $\|\mathcal{L}\varphi\|_{2}$, since for a fixed vertex on $l_{km}$ denoted with $v_{km}$, we would have had that
$|\mathcal{L}_{j}\varphi_{j}(v_{km})|=\left|\frac{1}{q+1}\varphi_{j}(v_{km+1})\right|$
while
$|\mathcal{L}\varphi(v_{km+1})|=\left|\frac{q}{q+1}\varphi_{j}(v_{km+1})+\frac{1}{q+1}\varphi(v_{km-1})\right|.$
This also means that our argument is valid if we remove the sphere $l_{1}$ from our sampling set.
An example for the case $k=3$ is given also in [12]: in this case, the finite sub-trees making up $S$ are just single nodes, so we can exploit a sharper threshold given by
$\|\varphi\|_{2}\leq\left(1+\frac{q}{(q+1)^{2}}\right)^{-1/2}\|\mathcal{L}\varphi\|_{2},\qquad\varphi\in L^{2}(l_{m}).$
By Lemma 2.7, $W$ would then be a viable sampling set for any $PW_{\omega}(G)$ space such that
$\omega<\left(1+\frac{q}{(q+1)^{2}}\right)^{1/2},$
which is of course a higher threshold then the one provided by our calculations.